OpenStudy (anonymous):

** AM I CORRECT?* Naomi uses a credit card with a 25.9% APR, compounded monthly, to pay for a hotel stay totaling $971.20. She can pay $400 per month on the card. What will the total cost of this purchase be? $1,009.36 $1,222.74 $1,180.56 $971.20 The answer I got was C

OpenStudy (anonymous):

What is apr?

OpenStudy (anonymous):

Annual payment rate?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Annual percentage rate - APR refers to the interest.

OpenStudy (anonymous):

Whoops

OpenStudy (anonymous):

Well what I did was 971.20*r - 400 = 592.16 592.16*r - 400 = 204.94 204.94*r = 209.37 Then I added that to 971.20

OpenStudy (anonymous):

Is the formula P(1+r/100)t used in this question?

OpenStudy (anonymous):

What did you use for "r"?

OpenStudy (anonymous):

0.259

OpenStudy (anonymous):

Rate

OpenStudy (anonymous):

Which was 25.9%

OpenStudy (anonymous):

Annual rate of interest

OpenStudy (anonymous):

Where is the t?

OpenStudy (anonymous):

Idk

OpenStudy (anonymous):

The tea is in the kitchen, would you like me to get you some?

OpenStudy (anonymous):

Bail?

OpenStudy (anonymous):

....

OpenStudy (anonymous):

Check this site out. It may help you: http://www.wikihow.com/Calculate-Annual-Percentage-Rate

OpenStudy (anonymous):

Good find

OpenStudy (anonymous):

:/

OpenStudy (anonymous):

I just want someone to check my answer ahhaha

OpenStudy (anonymous):

I'd love to help, but I don't know how to do it. Sorry. You may be able to find a calculator online that can check your answer.

OpenStudy (anonymous):

Neither can i sorry!

OpenStudy (anonymous):

I don't like using online calculators, but thank you guys for trying!

OpenStudy (agent0smith):

There is a formula for these questions, but I don't know it (i've only seen it a couple of times). You should have it in your notes or your book.

OpenStudy (agent0smith):

It's not P(1+r/100)t but it's similar, it accounts for the monthly payments.

OpenStudy (anonymous):

This is an online class, so there is no book or notes. I used the correct formula I think but I'm just making sure it's right by getting a second opinion :)

OpenStudy (agent0smith):

What was the formula you used? I'll remember it if i see it... I have a finite math book which has the formula in it.

OpenStudy (anonymous):

i = 0.259 -- Annual Interest j = i/12 -- Monthly Interest r = 1+j -- Monthly Accumulation Factor 971.20*r - 400 = 592.16 592.16*r - 400 = 204.94 204.94*r = 209.37

OpenStudy (anonymous):

1+n **

OpenStudy (anonymous):

I'll take my chances and say it's right, I'll tell you if it is or not in a sec!

OpenStudy (anonymous):

Poop, it was wrong

OpenStudy (agent0smith):

Are you sure you're using the right formula? If you post the actual formula, i might recognize it.

OpenStudy (agent0smith):

And what was your final answer? $1009.36?

OpenStudy (anonymous):

Yeah, the answer was probably 1009.36 I probably used the wrong formula.

OpenStudy (agent0smith):

Well... using your math i got 1009.36.

OpenStudy (agent0smith):

Add up the payments (i starred them): 971.20*r - ***400*** = 592.16 592.16*r - ***400*** = 204.94 204.94*r = ***209.37***

OpenStudy (agent0smith):

Those added together give the total amount paid.

OpenStudy (agent0smith):

The amounts on the right are just remaining balance, except the last one, that's the final payment. The other two payments are the 400s.